1. 12 HClO4 + P4O10 -> 4H3PO4 + 6 Cl2O7

given 5.48 mol HClO4, how many miles of H3PO4 could be produced (assuming abundant P4O10 is present)

  1. 8 CO + 17 H2 -> C8H18 + 8 H2O

assuming an unlimited supply of hydrogen gas (H2), how many miles of C8H18 could be produced with 26.79 miles of carbon monoxide?

  1. B2Br6 + 6HNO3 -> 2 B(NO3)3 + 6 HBr

given 276.2 g B2Br6, how many grams of HBr could be produced? assume you have more than enough HNO3



1
Expert's answer
2021-01-28T05:13:12-0500

Solution:

(1):

The balanced chemical equation:

12HClO4 + P4O10 → 4H3PO4 + 6Cl2O7

According to the chemical equation:

n(HClO4)/12 = n(H3PO4)/4

n(HClO4)/3 = n(H3PO4)

Hence,

n(H3PO4) = n(HClO4) / 3 = 5.48 mol / 3 = 1.82667 mol = 1.83 mol

n(H3PO4) = 1.83 mol

Answer (1): 1.83 moles of H3PO4 could be produced.


(2):

The balanced chemical equation:

8CO + 17H2 → C8H18 + 8H2O

According to the chemical equation:

n(CO)/8 = n(C8H18)

Hence,

n(C8H18) = n(CO) / 8 = 26.79 mol / 8 = 3.34875 mol = 3.35 mol

n(C8H18) = 3.35 mol

Answer (2): 3.35 moles of C8H18 could be produced.


(3):

The balanced chemical equation:

B2Br6 + 6HNO3 → 2B(NO3)3 + 6HBr

According to the chemical equation: n(B2Br6) = n(HBr)/6

or:

6 × m(B2Br6) / M(B2Br6) = m(HBr) / M(HBr)

The molar mass of B2Br6 is 501.046 g mol-1.

The molar mass of HBr is 80.912 g mol-1.

Hence,

m(HBr) = 6 × m(B2Br6) × M(HBr) / M(B2Br6)

m(HBr) = (6 × 276.2 g × 80.912 g mol-1) / (501.046 g mol-1) = 267.6 g

m(HBr) = 267.6 g

Answer (3): 267.6 grams of HBr could be produced.